OpenStudy (blackbird02):

How to solve this. 5^x-2(5^-x)+1=0

OpenStudy (anonymous):

use (ln)

OpenStudy (anonymous):

\[5^{x}-2(5^{-x})+1=0\] is this the equation

OpenStudy (blackbird02):

yes. then it would be: \[[5^x-2(\frac{ 1 }{ 5^x })+1=0][5^x]\] right?

OpenStudy (anonymous):

why there is 5^x in right side

OpenStudy (blackbird02):

to eliminate the denominator

OpenStudy (anonymous):

\[5^{x}-\frac{ 2 }{5^{x} }+1=0\]

OpenStudy (anonymous):

\[5^{x}\left[ 5^{x} -\frac{ 2 }{ 5^{x}}+1\right]=5^{x}(0)\] \[5^{2x}-2+5^{x}=0\] \[\ln (5^{2x})-\ln (2)+\ln )5^{5}=\ln (0)\] \[2x \ln (5)-\ln(2)+xln(5)=\ln(0)\] \[x(2\ln(5)+\ln(5)=\ln(0)+\ln(2)\] \[x=\frac{ \ln(0)+\ln(2) }{ 2\ln(5)+\ln(5) }\]

OpenStudy (michele_laino):

@blackbird02 try this change of variable: \[5^{x}=y\] where y is the new variable, substitute it in your equation, and you get a quadratic equation in y, the solve it for y

OpenStudy (blackbird02):

ok. thanks guys. i get it now